OpenStudy (anonymous):

How would I find a cubic polynomial with roots 1 and i? Thanks!

OpenStudy (y2o2):

(x-i)(x-1)² or (x-1)(x-i)²

OpenStudy (phi):

You need to know that complex roots come in complex conjugate pairs also, a cubic (3rd power) has 3 roots. for a quadratic with roots 2 and 3, the poylnomial is (x-2)(x-3)

OpenStudy (phi):

y2, not repeated i, but complex conjugate pairs.

OpenStudy (anonymous):

I am honestly trying to understand, but I am not sure how I would use that information to start the question. Is there a general process for this type of question?

OpenStudy (phi):

First, can you identify the 3 roots? Do you know what complex conjugate pair is?

OpenStudy (anonymous):

I'm not sure how to find the third root, but I realize that two of the roots are (1+0i) and (0+1i).

OpenStudy (phi):

no. The roots are 1, +i , and -i (the complex conjugate of a+bi is a-bi, here a=0, b=1)

OpenStudy (anonymous):

Oh, I see. Okay.

OpenStudy (phi):

Now we know the terms (x-i)(x- -i)(x-1)= p(x) or (x-i)(x+i)(x-1) = p(x) multiply out to get the full expression

OpenStudy (y2o2):

Oh, i thought he wanted the roots to be only +i and 1 i didn't know that -i is also included

OpenStudy (phi):

You always get imaginary roots in conjugate pairs (unless your polynomial has complex coeffs, which we typically do not have)

OpenStudy (anonymous):

Okay, thank you very much! I understand it now. :)

OpenStudy (y2o2):

Hmm, Ok, thanks for the information